Treatment

Health Benefit to Patient

(in $1000s)

Treatment Cost[1]

(in $1000s)

Mild

Severe

A

8

8

2

B

11

30

10

C

14

60

32

D

13

65

58

Four treatments exist to treat patients with a particular intestinal disorder. Patients can have either a “mild” or “severe” form of the disorder, which the physician can observe but the insurer cannot. One-quarter (25%) of the patients with this disorder have the “severe” form. The cost and health benefits (measured in terms of patients’ willingness-to-pay) associated with the four treatments are given in the table below.

  1. What is the efficient treatment for each type of patient? If treatments are efficient, what is the average treatment cost per treated patient? What is the average surplus created per treated patient?

.
Answer #1

The most efficient treatment will be the one with the largest difference between benefit and costs. So for mild patient, it is treatment A as the net benefit is 6, which is the highest. For severe, it is C as net benefit is 28, which is the highest.

Average treatment cost=0.25*32+0.75*2=9.5

Average surplus= 0.25*(60-32)+0.75*(8-2)=11.25